On April 1, 2018, Jiro Nozomi created a new travel agency, Adventure Travel. The following transactions occurred during the ca. Two-thirds (or $133) of one months insurance coverage has expired. b. At the end of the month, $600 of office supplies arPrepare the balance sheet at April 30, 2018. ADVENTURE TRAVEL Balance Sheet April 30, 2018 Cash $ Accounts receivable PrepaidSI need serious help with all of this I'm so confused
Answer #1

Part 1 & 2

Date

General journal

Debit

Credit

April 01

Cash

30000

Computer equipment

20000

J. Nozomi, capital

50000

April 02

Rent expense

1800

Cash

1800

April 03

Office supplies

1000

Cash

1000

April 10

Prepaid insurance

2400

Cash

2400

April 14

Salaries expense

1600

Cash

1600

April 24

Cash

8000

Commission earned

8000

April 28

Salaries expense

1600

Cash

1600

April 29

Repairs expense

350

Cash

350

April 30

Telephone expense

750

Cash

750

April 30

Dividends

1500

Cash

1500

Part 3

Transaction

General journal

Debit

Credit

a

Insurance expense

133

Prepaid insurance

133

b

Office supplies expense (1000-600)

400

Office supplies

400

c

Salaries expense

420

Salaries payable

420

d

Accounts receivable

1750

Commissions earned

1750

Part 4 a

Adventure Travel

Income Statement

For the month ending April 3018

Revenues:

Commissions earned

9750

Expenses:

Salaries expense

3620

Insurance expense

133

Rent expense

1800

Office supplies expense

400

Telephone expense

750

Repairs expense

350

Total expenses

7053

Net income

$2697

(8000+1750) = 9750

420+3200 = 3620

Part 4 b

Adventure Travel

Statement of retained earnings

For the month ending April 30, 2018

Retained earnings, April 1

0

Add: net income

2697

2697

Less: dividends

1500

Retained earnings, April 30

1197

Part 4 c

Adventure Travel

Balance Sheet

April 30, 2018

Assets

Cash

27000

Accounts receivable

1750

Office supplies

600

Prepaid insurance (2400-133)

2267

Total current assets

31617

computer

20000

Total assets

$51617

Liabilities and stockholders’ equity

Current liabilities

Salaries payable

420

Stockholder’s equity

J. Nozomi, capital

50000

Retained earnings

1197

Total stockholders’ equity

51197

Total liabilities and stockholders’ equity

$51617

Cash = 30000-1800-1000-2400-1600+8000-1600-350-750-1500 = 27000

Part 5

Date

General journal

Debit

Credit

April 30, 2018

Commissions earned (revenues)

9750

Income summary

9750

April 30, 2018

Income summary

7053

Salaries expense

3620

Insurance expense

133

Rent expense

1800

Office supplies expense

400

Telephone expense

750

Repairs expense

350

April 30, 2018

Income summary

2697

Retained earnings

2697

April 30, 2018

Retained earnings

1500

Dividend

1500